ltalloc.dtx

Johannes Braams  David Carlisle  Alan Jeffrey
Leslie Lamport  Frank Mittelbach  Chris Rowley

Rainer Schöpf

1995/10/25

1 Counters

This section deals with counter and other variable allocation.

1 ⟨∗2ekernel⟩

THE FOLLOWING ARE FROM PLAIN: \z@ : A zero dimen or number. It’s more efficient to write

              \parindent\z@ than \parindent 0pt.

\@ne : The number 1. \m@ne : The number -1. \tw@ : The number 2. \sixt@@n : The number 16. \@m : The number 1000. \@MM : The number 20000.

\@xxxii The constant 32.

2 \chardef\@xxxii=32
\@Mi \@Mii \@Miii \@miv Constants 1001–1004.
3 \mathchardef\@Mi=10001
4 \mathchardef\@Mii=10002
5 \mathchardef\@Miii=10003
6 \mathchardef\@Miv=10004

\@tempcnta \@tempcntb Scratch count registers used by L^{A}TEX kernel commands.

7 \newcount\@tempcnta
8 \newcount\@tempcntb

\if@tempswa General boolean switch used by L^{A}TEX kernel commands.

9 \newif\if@tempswa\@tempswatrue

\@tempdima \@tempdimb \@tempdimc Scratch dimen registers used by L^{A}TEX kernel commands.

10 \newdimen\@tempdima
11 \newdimen\@tempdimb
12 \newdimen\@tempdimc

\@tempboxa Scratch box register used by L^{A}TEX kernel commands.

             13 \newbox\@tempboxa

\@tempskipa \@tempskipb Scratch skip registers used by L^{A}TEX kernel commands.

             14 \newskip\@tempskipa
             15 \newskip\@tempskipb

\@temptokena Scratch token register used by L^{A}TEX kernel commands.

             16 \newtoks\@temptokena
\@flushglue Glue used for \right- & \leftskip = 0pt plus 1fil
             17 \newskip\@flushglue \@flushglue = 0pt plus 1fil

18 ⟨/2ekernel⟩